OpenStudy (anonymous):

Which placement of grouping symbols in the expression gives it a value of 22? 3 • 4 – 2 + 5 +7 A. 3 • [4 – (2 + 5) + 7] B. (3 • 4) – [2 + (5 + 7)] C. [(3 • 4) – 2] + (5 + 7) D. 3 • (4 – 2) + (5 + 7)

OpenStudy (md152727):

Have you started to answer this?

OpenStudy (anonymous):

Not yet.

OpenStudy (md152727):

[(3×4)−2]+(5+7)=[12−2]+12=10+12=22

OpenStudy (md152727):

So?

OpenStudy (anonymous):

B?

OpenStudy (md152727):

Nope, C

OpenStudy (anonymous):

AHHHH. Thank you!

OpenStudy (md152727):

The parenthesis matters, Your welcome